OpenStudy (anonymous):

Anyone good at linear algebra, msg me please! Need help in finding value of f to stabalize population matrix!! A =  [0.7 − f 0.2] [3 0 ]

OpenStudy (amistre64):

what does it mean to stabalize a population?

OpenStudy (amistre64):

cause I dont recall anything about stabalizing a matrix :)

OpenStudy (anonymous):

i believe that stabalizing means that if you start with a population, Xo = 80, and Yo = 30, if you keep multiplying the product of A and (Xo, Yo), then it willl converge towards a point .. if f is the right value, if not then the population will grow or die off.

OpenStudy (amistre64):

like a marchov chain

OpenStudy (anonymous):

uhm maybe?

OpenStudy (amistre64):

a markov chain takes the results of the the matrix and reapplies the matrix to it in a cycle

OpenStudy (amistre64):

Px = a ; Pa = b ; Pb = c ; ..... to stabilize i think would be to get to the point where Pn=n

OpenStudy (anonymous):

hmm ok. not sure how i would go about doing that.

OpenStudy (amistre64):

http://www.sosmath.com/matrix/markov/markov.html this might help a little, im reading it now to refresh the brain

OpenStudy (amistre64):

lol, eugene value of 1

OpenStudy (anonymous):

ahahh ugh i suck.

OpenStudy (amistre64):

[0.7 − f 0.2] [x] = [x] [3 0 ] [y] [y] this creates a system of equations such that x(.7-f) + .2y = x 3x + 0y = y

OpenStudy (amistre64):

x(.7-f) + .2y = x .7x -fx -x +.2y = 0 3x + 0y = y 3x -y = 0 (.7-1-f)x +.2y = 0 3x - y = 0

OpenStudy (amistre64):

-.3-f .2 3 -1 3 -1 -.3-f .2 1 -1/3 -1 .2/(.3-f) 1 -1/3 0 .2/(.3-f) - 1/3 .2/(.3-f) - 1/3 = 1 .2/(.3-f) = 4/3 1/(.3-f) = 4/.6 .3-f = .6/4 -f = .6/4 - .3 f = .3- .6/4 maybe

OpenStudy (amistre64):

with any luck, that means that when f - .15 its stabalized

OpenStudy (amistre64):

when f = .15 ....

OpenStudy (anonymous):

ok, i got f = .5 but i did it a completely different way, yours looks a little more right ..

OpenStudy (amistre64):

lets try them out, cause mine could have errors in it :)

OpenStudy (anonymous):

ok, how are we supposed to test the value? ahah.

OpenStudy (anonymous):

just put (80, 30) through a munch of times i'm guessing.

OpenStudy (amistre64):

hmm, yeah that should work :) does the vector need to equal 100% tho? i cant recall

OpenStudy (anonymous):

i have no idea to be honest, i dont even know what i'm looking for.

OpenStudy (amistre64):

i remember someting :) maybe

OpenStudy (amistre64):

the eugene vector associated with a eugen value of 1 will give us Px = x

OpenStudy (amistre64):

so we want the determinant of this given matrix to be zero

OpenStudy (amistre64):

since 1 times anything is itself; we dont need to modify anything

OpenStudy (amistre64):

well, -1 the diags

OpenStudy (anonymous):

ok! thats how i got f = .5 for my answer, but as i'm running it through it, the pop'n doesn't seem to be stabalizing

OpenStudy (amistre64):

[0.7−f-1 0.2 ] [ 3 0-1 ] thats what I did too in another fashion lol -(-0.3-f) - (0.6) = 0 0.3+f = 0.6 f = 0.3

OpenStudy (amistre64):

0.7 - 0.3 = 0.4

OpenStudy (amistre64):

you sure you got f=.5?

OpenStudy (anonymous):

i probably did it wrong, but i'm doing trial and analysis, and .4 would probably work! i'm going to try it!!!

OpenStudy (anonymous):

ok so i think it works, stabalizing at (54, 161)

OpenStudy (amistre64):

yay!! :)

OpenStudy (anonymous):

does that seem right to you?!

OpenStudy (amistre64):

dunno, i aint got a program to simulate it with :/

OpenStudy (amistre64):

we could try crossing lines is my thought about it, but then lines are not "feedback" loops

OpenStudy (anonymous):

hmmm, i dont even know how to do that haha, i suck. can i ask you one more question that is another part of this question? If initially there are no fish in our lake, and we wish to stock it with adult fish (and no young fish) so that the total number of fish is 10, 000 when it hits the stable level (counting both adult and young fish together, and with the f from above), how many adult fish should we introduce into the lake to make this happen?

OpenStudy (amistre64):

i would have to know the information that the matrix was made with in order to know how to use this new information

OpenStudy (anonymous):

the matrix is just the new one that we made i think where it stabalizes with .4 .2 3 0

OpenStudy (amistre64):

right, and we have to use a vector that resembles [adult,young] or is it [young, adult] to run a scenario with

OpenStudy (amistre64):

the matrix itself does not exist in a vaccuum; it was created by some information that tells me how to use this new information; otherwise I have no direction to go with

OpenStudy (anonymous):

(adult, young), i'm not sure then .. would it be the first matrix we were working with without the f?

OpenStudy (anonymous):

no it wouldn't be that matrix becuase the pop'n doesnt stabalize in that matrix.

OpenStudy (amistre64):

is the left column vector represent adult information or young information?

OpenStudy (amistre64):

the numbers are evidently a yearly cycle of how many live and die

OpenStudy (amistre64):